# Dynamic asymptotic dimension growth for group actions and groupoids

### v1

@@ 5. Partitions of unity
In this section, we develop a critical tool that is not only a partition of unity for slow dynamic asymptotic dimension growth, but also provides good properties and estimates. Therefore, this tool plays a key role in proving the amenability of groupoids in the next section. Throughout this section, $\mathcal{G}$ is an étale groupoid, and $\ell$ is a coarse continuous length function on $\mathcal{G}$.
We are now ready to formulate our generalization of [[GWY17], Proposition 7.1].
@@
To clearly understand the above proposition, we look at the case of group actions on compact spaces.
@@
To obtain the main proposition of this section, we need several lemmas. The first step is to show the following lemma, which plays an important auxiliary role in the proof of subsequent lemmas and theorems.
@@
If the unit space of an étale groupoid is compact and Hausdorff, then it is normal, which yields an excellent result as follows.
@@
We divide our proof into four steps. First, we show that for any open subset $V$ of $\mathcal{G}^{(0)}$ and any $x\in V$, there exists an open neighborhood $U_{x}$ of $x$ such that $\overline{U_{x}}\subseteq V$. Note that $\mathcal{G}^{(0)}$ is Hausdorff, thus the singleton set $\{x\}$ is closed. Also, $\mathcal{G}^{(0)}$ is compact, then it is normal. Observe that the boundary point $\partial V$ is closed, and $\{x\}\cap\partial V\subseteq V\cap\partial V=\varnothing$ as $V$ is open. Hence by normality of $\mathcal{G}^{(0)}$, there exist an open neighborhood $U_{x}$ of $x$ and open set $V^{\prime}$ containing $\partial V$ such that
@@
We may assume that $U_{x}\subseteq V$, otherwise we replace $U_{x}$ by $U_{x}\cap V$. It suffices to show that $\overline{U_{x}}\subseteq V$. If this does not happen, we may find $y\in\partial U_{x}\subseteq\overline{V}$ such that $y\notin V$, which implies $y\in\partial V\subseteq V^{\prime}$. It follows that $V^{\prime}$ is an open neighborhood of $y$, and that $U_{x}\cap V^{\prime}\neq\varnothing$ as $y\in\partial U_{x}$, contradiction. Hence, $\overline{U_{x}}\subseteq V$ as required.
@@
Another step is to construct subsets $V_{i}^{\prime}$ and $\mathcal{U}_{i}$ of $\mathcal{G}^{(0)}$ such that $U_{y}\subseteq V_{i}^{\prime}$ when $y\in\partial V_{0}\cap V_{i}$ for some $i\in\{1,\cdots,n\}$, and such that $V_{i}^{\prime}\subseteq\mathcal{U}_{i}\subseteq\overline{\mathcal{U}_{i}}\subseteq V_{i}$. From the second step, we see that $U_{y}\subseteq V_{i}$ and $U_{y}\cap\widetilde{U_{z}}=\varnothing$, $\forall z\in\partial V_{i}$. For fixed $i$ and $z\in\partial V_{i}$, $z\notin V_{i}$ implies that there is at least one $l\in\{0,\cdots,i-1,i+1,\cdots,n\}$ such that $z\in V_{l}$. Therefore, there exists an open neighborhood $U_{z}^{\prime}$ of $z$ such that $U_{z}^{\prime}\subseteq V_{l}$ as $V_{l}$ is open. Similar to the second step, for any $h\in\partial V_{l}$, there exist open neighborhoods $\widetilde{U_{h}}$ of $h$ and $U_{z}^{\prime\prime}$ of $z$ such that $\widetilde{U_{h}}\cap U_{z}^{\prime\prime}=\varnothing$. Replacing $U_{z}^{\prime}$ by $U_{z}^{\prime}\cap U_{z}^{\prime\prime}$, we have $\widetilde{U_{h}}\cap U_{z}^{\prime}=\varnothing$ and $U_{z}^{\prime}\subseteq V_{l}$. Set $U_{z}=\widetilde{U_{z}}\cap U_{z}^{\prime}$, then $U_{z}\subseteq V_{l}$, $U_{z}\cap U_{y}=\varnothing$ and $U_{z}\cap\widetilde{U_{h}}=\varnothing$. Let $V_{i}^{\prime}=\overline{V_{i}}\backslash\bigcup\limits_{z\in\partial V_{i}}U_{z}$, by $U_{z}\cap U_{y}=\varnothing$, then we get the desired inclusion $U_{y}\subseteq V_{i}^{\prime}$. Repeating the second step, we may obtain an open subset $\mathcal{U}_{i}$ such that $V_{i}^{\prime}\subseteq\mathcal{U}_{i}\subseteq\overline{\mathcal{U}_{i}}\subseteq V_{i}$.
@@
With the help of the preceding two lemmas, we can now acquire the following key lemma, which is an adaption of [[GWY17], Lemma 7.4].
@@
Since $\overline{K}$ is compact, we see that $\overline{K}^{n}$ is compact by Lemma 4.8. Also, $\overline{K}^{n}\subseteq\overline{K}^{n+1}$ for each $n\in\mathbb{N}$ as long as we let $g_{n+1}=r(g_{n})\in\overline{K}$. Observe that for any $z\in\overline{K}^{N(R)}$, $\ell(z)\leq N(R)R<N(R)R+1$, thus $\overline{K}^{N(R)}$ is contained in $K^{\prime}=\{z\in\mathcal{G}\mid\ell(z)<N(R)R+1\}$, which is an open precompact symmetric subset of $\mathcal{G}$ containing $\mathcal{G}^{(0)}$. According to Lemma 5.4, there are open precompact subsets $V_{0},\cdots,V_{f(N(R)R+1)}$ of $\mathcal{G}^{(0)}$ that cover $\mathcal{G}^{(0)}$ such that for each $i$, the set $\{g\in K^{\prime}\mid s(g),r(g)\in V_{i}\}$ generates an open precompact subgroupoid of $\mathcal{G}$.
@@
So far, we have built the tools required for Proposition 5.1, and now we can start proving it.
@@
Step 1. In the first step, we build a desired open cover of $\mathcal{G}^{(0)}$. Without loss of generality, we may assume that $\mathcal{G}$ has dynamic asymptotic dimension function $f$. Given $R>0$, the open precompact subset $K=\{z\in\mathcal{G}\mid\ell(z)<R\}$ of $\mathcal{G}$, and $\varepsilon>0$. Clearly, $K=K^{-1}$, $s(K)\cup r(K)=\mathcal{G}^{(0)}\subseteq K$. Let $p=[\frac{\alpha}{1-\alpha}]+1$, and let $N(R)=[(R+1)^{c}]$ for some constant $c$ in Lemma 5.6, such that
@@
such constant $c$ exists and depends on given parameters $\alpha,R,\varepsilon$. Let $\{U_{i}^{(n)}\mid i\in\{0,\cdots,f(N(R)R+1)\},n\in\{0,\cdots,N(R)\}\}$ have the properties in Lemma 5.6. For each $i$, let $U_{i}=U_{i}^{(N(R))}$. By Lemma 5.6 (i)(ii), $\{U_{0}^{(0)},\cdots,U_{f(N(R)R+1)}^{(0)}\}$ covers $\mathcal{G}^{(0)}$ and $U_{i}^{(0)}\subseteq U_{i}^{(N(R))}=U_{i}$, thus the family $\{U_{0},\cdots,U_{f(N(R)R+1)}\}$ covers $\mathcal{G}^{(0)}$. By Lemma 5.6 (iii), we have that $s(K\cap r^{-1}(U_{i}^{(0)}))\subseteq U_{i}^{(1)}\subseteq U_{i}$ for some $i$, which implies that for any $x\in\mathcal{G}^{(0)}$, $s(K\cap r^{-1}(x))\subseteq U_{i}$ for some $U_{i}$. By Lemma 5.6 (iv), we see that $\{g\in K\mid s(g),r(g)\in U_{i}\}$ generates an open precompact subgroupoid of $\mathcal{G}$.
Step 2. The second step is to construct a partition of unity. By Lemma 5.6 (ii), for each $i\in\{0,\cdots,f(N(R)R+1)\}$, and each $n\in\{1,\cdots,N(R)\}$, $\overline{U_{i}^{(n-1)}}\subseteq U_{i}^{(n)}$. According to Urysohn’s lemma, there exist continuous functions
@@
such that $\psi_{i}^{(n)}(x)=1$, $\forall x\in U_{i}^{(n-1)}$; $\psi^{(n)}_{i}(x)=0$, $\forall x\in\mathcal{G}^{(0)}\backslash U_{i}^{(n)}$.
@@
then supp ($\textnormal{supp }(\psi_{i})\subseteq U_{i}$ and $\psi_{i}(x)\leq 1,\forall x\in\mathcal{G}^{(0)}$. Define
@@
which are continuous functions and supp ($\textnormal{supp }(\phi_{i})\subseteq U_{i}$. By Lemma 5.6 (i), for any $x\in\mathcal{G}^{(0)}$, there exists at least one $j\in\{0,\cdots,f(N(R)R+1)\}$ such that $x\in U_{j}^{(0)}$, thus $\psi_{j}(x)=1$, so $\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(x)\right)^{\frac{1}{p}}\geq 1$, therefore
@@
Since $g\in K$ and $r(g)\in U_{j}^{(M)}$, $g\in K\cap r^{-1}(U_{j}^{(M)})$. By Lemma 5.6 (ii), we see that $s(g)\in s(K\cap r^{-1}(U_{j}^{(M)}))\subseteq U_{j}^{(M+1)}$. Note that $r(g)\notin U_{j}^{(M-1)}\supseteq s(K\cap r^{-1}(U_{j}^{(M-2)}))$, thus $s(g^{-1})=r(g)\notin s(K\cap r^{-1}(U_{j}^{(M-2)}))$. We see that $g^{-1}\in K$ as $K=K^{-1}$, thus $g^{-1}\notin r^{-1}(U_{j}^{(M-2)})$, which implies that $s(g)=r(g^{-1})\notin U_{j}^{(M-2)}$. Therefore $s(g)\in U_{j}^{(M+1)}\backslash U_{j}^{(M-2)}$, and
@@
Combining equation 5.2 and equation 5.3, we have that
@@
Combining equation 5.4 and equation 5.5, we have
@@
On the other hand, using equations 5.4 and 5.5 again, we obtain
@@
where the third inequality comes from the Minkowski inequality
@@
According to equations 5, 5.7, and 5, we have
@@
Finally, from the equation 5.1, we obtain
@@
Given $10<\alpha<$ 1, $R>0$ and $\varepsilon>0$. To see the existence of $c$, we consider such example: $f(x)=[x^{\alpha}]$ , $p=[\frac{\alpha}{1-\alpha}]+1$. Through precise calculations, we found that if the constant $c$ satisfies the following inequality
@@
In fact, the finite dynamic asymptotic dimension also deduces almost invariant partitions of unity without the restriction of the compactness of the unit space, which is a suitable adaption of [Proposition 7.1, [GWY17]]. Hence, we get a stronger result as follows.
@@
    There exists a collection of continuous functions $\{\phi_{i}:\mathcal{G}^{(0)}\rightarrow[0,1]\}_{i=0}^{d}$ such that the support of each $\phi_{i}$ is contained in $U_{i}$, $\sum\limits_{i=0}^{d}\phi_{i}(x)\leq 1$ for all $x\in\mathcal{G}^{(0)}$, and $\sum\limits_{i=0}^{d}\phi_{i}(x)=1$ for all $x\in r(K)\cup s(K)$, and such that for any $g\in K$ and each $i$,
@@ 6. Amenability
Our motivation for discussing the relationship between the dynamic asymptotic dimension growth of a groupoid and the amenability is fourfold. First, in the work of [Tu99], Tu showed that the Baum-Connes conjecture holds for foliations whose holonomy groupoid is Hausdorff and amenable, particularly for amenable groupoids. Second, for a countable discrete group $\Gamma$ with property A, Higson showed that the Novikov conjecture holds [Hig00]. If such $\Gamma$ has bounded geometry, then property A is equivalent to the amenability of its associated coarse groupoid $G(\Gamma)$ [STY02]. Therefore, if a countable discrete group with bounded geometry has the amenable coarse groupoid, then the Novikov conjecture holds. Third, according to Corollary 4.15, we see that the subexponential dynamic asymptotic dimension growth of a coarse groupoid implies the amenability. Finally, analogous to the fact that finite dynamical complexity of a groupoid implies the amenability [GWY22], we prove that every locally compact Hausdorff étale groupoid $\mathcal{G}$ with finite dynamic asymptotic dimension is amenable by using the optimized Proposition 5.8. More generally, applying Proposition 5.1 in the previous section, we directly prove that every $\sigma$-compact, locally compact Hausdorff étale groupoid $\mathcal{G}$ with compact unit space having dynamic asymptotic dimension growth at most $x^{\alpha}$ $(0<\alpha<1)$ is amenable.
A reference [AR00] provides a standard description of an amenable groupoid and another reference [BO08] depicts the étale case. The following equivariant definition is a slight variant of [Definition 5.6.13, [BO08]], which is used in the proofs of two theorems of this section.
Definition 6.1 ([GWY22]).
@@
The following lemma is a vital tool in the proofs of two theorems.
Lemma 6.2 ([GWY22]).
@@
The following theorem is inspired in part by Guentner,Willett, and Yu’s work [GWY22] on dynamical complexity and controlled operator $K$-theory, and their work [GWY17] on finite dynamic asymptotic dimension. With the help of the nuclear dimension for the reduced groupoid $C^{*}$-algebra, they showed that a free locally compact Hausdorff étale groupoid with finite dynamic asymptotic dimension is amenable [[GWY17], Corollary 8.25]. Unlike their proof method, we adopt a direct method analogous to [[GWY22], Theorem A.9]; hence, we strengthen their result by removing the freeness assumption on groupoids.

### v2

@@ 5. Partitions of unity
The goal of this section is to develop a critical tool called a generalized partition of unity. This tool serves not only as a partition of unity for groupoids with slow dynamic asymptotic dimension growth, but also provides almost invariant properties and beneficial estimates. Furthermore, this tool is essential for demonstrating the amenability of groupoids in the following section.
The key techniques needed in the proof of the amenability of groupoids in the next section are summarized in the following proposition. We are now ready to present our key technical proposition, which generalizes [[GWY17], Proposition 7.1].
@@
To understand the above proposition more intuitively, we look at the case of group actions on compact spaces.
@@
To obtain the main proposition of this section, we need a couple of lemmas. The first step is to show the following lemma, which plays an important auxiliary role in the proof of subsequent lemmas and theorems.
@@
If the unit space of an étale groupoid is compact and Hausdorff, then it is normal, leading to an excellent result as follows.
@@
We divide our proof into four steps. First, we show that for any open subset $V$ of $\mathcal{G}^{(0)}$ and any $x\in V$, there exists an open neighborhood $U_{x}$ of $x$ such that $\overline{U_{x}}\subseteq V$. Note that $\mathcal{G}^{(0)}$ is Hausdorff, thus the singleton set $\{x\}$ is closed. Additionally, since $\mathcal{G}^{(0)}$ is compact, it is also normal. Observe that the boundary $\partial V$ is closed, and $\{x\}\cap\partial V\subseteq V\cap\partial V=\varnothing$ as $V$ is open. Hence by normality of $\mathcal{G}^{(0)}$, there exist an open neighborhood $U_{x}$ of $x$ and open set $V^{\prime}$ containing $\partial V$ such that
@@
We may assume that $U_{x}\subseteq V$, otherwise we can replace $U_{x}$ with $U_{x}\cap V$. It suffices to show that $\overline{U_{x}}\subseteq V$. If this does not happen, we may find a point $y\in\partial U_{x}\subseteq\overline{V}$ such that $y\notin V$, which implies $y\in\partial V\subseteq V^{\prime}$. It follows that $V^{\prime}$ is an open neighborhood of $y$, and that $U_{x}\cap V^{\prime}\neq\varnothing$ as $y\in\partial U_{x}$, contradiction. Hence, we conclude that $\overline{U_{x}}\subseteq V$, as required.
@@
Another step is to construct subsets $V_{i}^{\prime}$ and $\mathcal{U}_{i}$ of $\mathcal{G}^{(0)}$ such that $U_{y}\subseteq V_{i}^{\prime}$ when $y\in\partial V_{0}\cap V_{i}$ for some $i\in\{1,\cdots,n\}$, and such that $V_{i}^{\prime}\subseteq\mathcal{U}_{i}\subseteq\overline{\mathcal{U}_{i}}\subseteq V_{i}$. From the second step, we see that $U_{y}\subseteq V_{i}$ and $U_{y}\cap\widetilde{U_{z}}=\varnothing$, $\forall z\in\partial V_{i}$. For fixed $i$ and $z\in\partial V_{i}$, $z\notin V_{i}$ implies that there is at least one $l\in\{0,\cdots,i-1,i+1,\cdots,n\}$ such that $z\in V_{l}$. Therefore, there exists an open neighborhood $U_{z}^{\prime}$ of $z$ such that $U_{z}^{\prime}\subseteq V_{l}$ as $V_{l}$ is open. Similar to the second step, for any $h\in\partial V_{l}$, there exist open neighborhoods $\widetilde{U_{h}}$ of $h$ and $U_{z}^{\prime\prime}$ of $z$ such that $\widetilde{U_{h}}\cap U_{z}^{\prime\prime}=\varnothing$. Replacing $U_{z}^{\prime}$ by $U_{z}^{\prime}\cap U_{z}^{\prime\prime}$, we have $\widetilde{U_{h}}\cap U_{z}^{\prime}=\varnothing$ and $U_{z}^{\prime}\subseteq V_{l}$. Set $U_{z}=\widetilde{U_{z}}\cap U_{z}^{\prime}$, then $U_{z}\subseteq V_{l}$, $U_{z}\cap U_{y}=\varnothing$ and $U_{z}\cap\widetilde{U_{h}}=\varnothing$. Let $V_{i}^{\prime}=\overline{V_{i}}\backslash\bigcup\limits_{z\in\partial V_{i}}U_{z}$. Since $U_{z}\cap U_{y}=\varnothing$, we get the desired inclusion $U_{y}\subseteq V_{i}^{\prime}$. Repeating the second step, we may obtain an open subset $\mathcal{U}_{i}$ such that $V_{i}^{\prime}\subseteq\mathcal{U}_{i}\subseteq\overline{\mathcal{U}_{i}}\subseteq V_{i}$.
@@
The following lemma constructs a nested family of open precompact subsets of $\mathcal{G}^{(0)}$ that satisfy several favorable properties. This is the main technical ingredient needed to prove the main proposition.
@@
Since $\overline{K}$ is compact, it follows from Lemma 4.11 that $\overline{K}^{n}$ is also compact. Furthermore, we have $\overline{K}^{n}\subseteq\overline{K}^{n+1}$ for each $n\in\mathbb{N}$, provided we set $g_{n+1}=r(g_{n})\in\overline{K}$. For any $z\in\overline{K}^{N(R)}$, we can observe that $\ell(z)\leq N(R)R<N(R)R+1$, thus $\overline{K}^{N(R)}$ is contained in $K^{\prime}=\{z\in\mathcal{G}\mid\ell(z)<N(R)R+1\}$, which is an open precompact symmetric subset of $\mathcal{G}$ containing $\mathcal{G}^{(0)}$. According to Lemma 5.4, there exist open precompact subsets $V_{0},\cdots,V_{f(N(R)R+1)}$ of $\mathcal{G}^{(0)}$ that cover $\mathcal{G}^{(0)}$ such that for each $i$, the set $\{g\in K^{\prime}\mid s(g),r(g)\in V_{i}\}$ generates an open precompact subgroupoid of $\mathcal{G}$.
@@
We have established the necessary lemmas for Proposition 5.1, enabling us to commence its proof.
@@
Step 1. In the first step, we build a desired open cover of $\mathcal{G}^{(0)}$. Given $R>0$, the open precompact subset $K=\{z\in\mathcal{G}\mid\ell(z)<R\}$ of $\mathcal{G}$, and $\varepsilon>0$. Clearly, $K=K^{-1}$, $s(K)\cup r(K)=\mathcal{G}^{(0)}\subseteq K$. Let $p=[\frac{\alpha}{1-\alpha}]+1$, and let $N(R)=[(R+1)^{c}]$ for some constant $c$ in Lemma 5.6, such that
@@
such constant $c$ exists and depends on given parameters $\alpha,R,\varepsilon$. Let $\{U_{i}^{(n)}\mid i\in\{0,\cdots,f(N(R)R+1)\},n\in\{0,\cdots,N(R)\}\}$ have the properties in Lemma 5.6. For each $i$, let $U_{i}=U_{i}^{(N(R))}$. By Lemma 5.6 (1)(2), $\{U_{0}^{(0)},\cdots,U_{f(N(R)R+1)}^{(0)}\}$ covers $\mathcal{G}^{(0)}$ and $U_{i}^{(0)}\subseteq U_{i}^{(N(R))}=U_{i}$, thus the family $\{U_{0},\cdots,U_{f(N(R)R+1)}\}$ covers $\mathcal{G}^{(0)}$. By Lemma 5.6 (3), we have that $s(K\cap r^{-1}(U_{i}^{(0)}))\subseteq U_{i}^{(1)}\subseteq U_{i}$ for some $i$, which implies that for any $x\in\mathcal{G}^{(0)}$, $s(K\cap r^{-1}(x))\subseteq U_{i}$ for some $U_{i}$. By Lemma 5.6 (4), we see that $\{g\in K\mid s(g),r(g)\in U_{i}\}$ generates an open precompact subgroupoid of $\mathcal{G}$.
Step 2. The second step is to construct a partition of unity. By Lemma 5.6 (2), for each $i\in\{0,\cdots,f(N(R)R+1)\}$, and each $n\in\{1,\cdots,N(R)\}$, $\overline{U_{i}^{(n-1)}}\subseteq U_{i}^{(n)}$. According to Urysohn’s lemma, there exist continuous functions
@@
such that $\psi_{i}^{(n)}(x)=1$, for all $x\in U_{i}^{(n-1)}$; $\psi^{(n)}_{i}(x)=0$, for all $x\in\mathcal{G}^{(0)}\backslash U_{i}^{(n)}$.
@@
then supp ($\textnormal{supp }(\psi_{i})\subseteq U_{i}$ and $\psi_{i}(x)\leq 1,\text{ for all }x\in\mathcal{G}^{(0)}$. Define
@@
which are continuous functions and supp ($\textnormal{supp }(\phi_{i})\subseteq U_{i}$. By Lemma 5.6 (1), for any $x\in\mathcal{G}^{(0)}$, there exists at least one $j\in\{0,\cdots,f(N(R)R+1)\}$ such that $x\in U_{j}^{(0)}$, thus $\psi_{j}(x)=1$, so $\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(x)\right)^{\frac{1}{p}}\geq 1$, therefore
@@
Since $g\in K$ and $r(g)\in U_{j}^{(M)}$, $g\in K\cap r^{-1}(U_{j}^{(M)})$. By Lemma 5.6 (2), we see that $s(g)\in s(K\cap r^{-1}(U_{j}^{(M)}))\subseteq U_{j}^{(M+1)}$. Note that $r(g)\notin U_{j}^{(M-1)}\supseteq s(K\cap r^{-1}(U_{j}^{(M-2)}))$, thus $s(g^{-1})=r(g)\notin s(K\cap r^{-1}(U_{j}^{(M-2)}))$. We see that $g^{-1}\in K$ as $K=K^{-1}$, thus $g^{-1}\notin r^{-1}(U_{j}^{(M-2)})$, which implies that $s(g)=r(g^{-1})\notin U_{j}^{(M-2)}$. Therefore $s(g)\in U_{j}^{(M+1)}\backslash U_{j}^{(M-2)}$, and
@@
Combining equation (5.2) and equation (5.3), we have that
@@
Combining equation (5.4) and equation (5.5), we have
@@
On the other hand, using equations (5.4) and (5.5) again, we obtain
@@
where the second inequality comes from the Minkowski inequality
@@
According to equations (5), (5.7), and (5), we have
@@
Finally, from the equation (5.1), we obtain
@@
Let $10<\alpha<$ 1, $R>0$ and $\varepsilon>0$. We consider the function $f(x)=[x^{\alpha}]$, $p=[\frac{\alpha}{1-\alpha}]+1$ to examine the existence of the constant $c$. Through precise calculations, we found that if the constant $c$ satisfies the following inequality:
@@
After thoroughly examining the proof of Proposition 7.1 in [GWY17], we conclude that finite dynamic asymptotic dimension leads to an almost invariant partition of unity without the restriction of the compactness of the unit space. Consequently, we can present a stronger result as follows:
@@
    There exists a collection of continuous functions $\{\phi_{i}:\mathcal{G}^{(0)}\rightarrow[0,1]\}_{i=0}^{d}$ such that supp ($\textnormal{supp }(\phi_{i})\subseteq U_{i}$ for each $i$, $\sum\limits_{i=0}^{d}\phi_{i}(x)\leq 1$ for all $x\in\mathcal{G}^{(0)}$, and $\sum\limits_{i=0}^{d}\phi_{i}(x)=1$ for all $x\in r(K)\cup s(K)$, and such that for any $g\in K$ and each $i$,
@@ 6. Amenability
Our motivations for exploring the relationship between the dynamic asymptotic dimension growth of groupoids and amenability are threefold. Firstly, Guentner, Willett, and Yu showed that the finite dynamic asymptotic dimension of a free, locally compact Hausdorff étale groupoid implies amenability [GWY17]. Secondly, they demonstrated that the finite dynamical complexity of a locally compact Hausdorff étale groupoid implies amenability [GWY24]. It is worth noting that the finite dynamic asymptotic dimension leads to finite dynamical complexity when the dynamic asymptotic dimension is either zero or one. Lastly, according to Corollary 4.19 in the previous section, we see that for a discrete metric space with bounded geometry $X$, subexponential dynamic asymptotic dimension growth of the coarse groupoid $G(X)$ implies amenability.
Based on the above methods and results, we can also find sufficient conditions for the amenability of groupoids. By combining the ideas in [GWY24] with the optimized Proposition 5.8, we can establish the theorem that every locally compact Hausdorff étale groupoid $\mathcal{G}$ with finite dynamic asymptotic dimension is amenable, strengthening the result of [GWY17]. Furthermore, using a similar approach and Proposition 5.1, we can directly demonstrate that every $\sigma$-compact, locally compact Hausdorff étale groupoid $\mathcal{G}$ with compact unit space, and with dynamic asymptotic dimension growth at most $x^{\alpha}$ $(0<\alpha<1)$, is also amenable, thus generalizing Corollary 4.19 in the case of slow growth.
A reference [AR00] provides a standard description of an amenable groupoid, while another reference [BO08] addresses the étale case. The following equivariant definition is a slight variation of Definition 5.6.13 from [BO08] and is used in the proofs of two theorems in this section.
Definition 6.1 ([GWY24]).
@@
The following lemma demonstrates that an open precompact subgroupoid of an étale groupoid is amenable. This fact is crucial in the proofs of two theorems in this section.
Lemma 6.2 ([GWY24]).
@@
The following theorem is inspired in part by the work of Guentner, Willett, and Yu showing that finite dynamical complexity of a groupoid implies amenability [GWY24], and their earlier research on finite dynamic asymptotic dimension of étale groupoids [GWY17], as well as the fact that finite dynamic asymptotic dimension leads to finite dynamical complexity when the dynamic asymptotic dimension is either zero or one. They demonstrated that a free, locally compact, Hausdorff étale groupoid with finite dynamic asymptotic dimension is amenable by using the nuclear dimension for the reduced groupoid $C^{*}$-algebra (see [GWY17], Corollary 8.25). In contrast to their proof method in [GWY17], we adopt a direct approach similar to that in [[GWY24], Theorem A.9]. This allows us to strengthen their result [[GWY17], Corollary 8.25] by removing the freeness assumption for groupoids.